OpenStudy (anonymous):

What is the value of x so the rectangles have the same area. A. 12 & X B. X-2 & 16

OpenStudy (anonymous):

12x=16(x-2) 12x=16x-32 -16x=-16x-32 -------------- -4x/-4=-32/-4 x= 8